# LlamaFactory: Unified Efficient Fine-Tuning of 100+ Language Models

### v3

@@ Abstract
Efficient fine-tuning is vital for adapting large language models (LLMs) to downstream tasks. However, it requires non-trivial efforts to implement these methods on different models. We present LlamaFactory, a unified framework that integrates a suite of cutting-edge efficient training methods. It provides a solution for flexibly customizing the fine-tuning of 100+ LLMs without the need for coding through the built-in web UI LlamaBoard. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks. It has been released at https://github.com/hiyouga/LLaMA-Factory and received over 24,000 stars and 3,000 forks.
@@ 1 Introduction
LlamaFactory consists of three main modules: Model Loader, Data Worker and Trainer. We minimize the dependencies of these modules on specific models and datasets, allowing the framework to flexibly scale to hundreds of models and datasets. Concretely, we first establish a model registry where the Model Loader can precisely attach adapters to the pre-trained models by identifying exact layers. Then we develop a data description specification that allows the Data Worker to gather datasets by aligning corresponding columns. Furthermore, we provide plug-and-play implementations of efficient fine-tuning methods that enable the Trainer to active by replacing default ones. Our design allows these modules to be reused across different training approaches, significantly reducing integration cost of new methods.
@@
LlamaFactory is open-sourced under the Apache-2.0 license. It has already garnered over 24,000 stars and 3,000 forks on the GitHub, and hundreds of open-source models have been built upon LlamaFactory on the Hugging Face Hub [^1] https://huggingface.co/models?other=llama-factory. For example, Truong et al. (2024) build GemSUra-7B based on LlamaFactory, revealing the cross-lingual abilities of Gemma (Mesnard et al., 2024). Furthermore, dozens of studies have utilized our framework to explore LLMs (Wang et al., 2023a; Yu et al., 2023; Bhardwaj et al., 2024).
@@ 4.1 Model Loader
This section initially presents the four components in Model Loader: model initialization, model patching, model quantization, and adapter attaching, followed by a description of our approach of adapting to a wide range of devices by handling the parameter float-point precision during fine-tuning.
@@ 4.4 Utilities
During inference time, we reuses the chat template from the Data Worker to build the model inputs. We offer support for sampling the model outputs using Transformers (Wolf et al., 2020) and vLLM (Kwon et al., 2023), both of which support stream decoding. Additionally, we implement an OpenAI-style API that utilizes the asynchronous LLM engine and paged attention of vLLM to provide high-throughput concurrent inference services, facilitating the deployment of fine-tuned LLMs into various applications.
@@ 4.5 LlamaBoard: A Unified Interface for LlamaFactory
LlamaBoard allows us to customize the fine-tuning arguments through interaction with the web interface. We provide default values for a majority of arguments that are recommended for most users, simplifying the configuration process. Moreover, users can preview the datasets on the web UI to check their correctness.
@@
LlamaBoard supports calculating the text similarity scores on the datasets to automatically evaluate the model or performing human evaluation by chatting with the model.

### v4

@@ Abstract
Efficient fine-tuning is vital for adapting large language models (LLMs) to downstream tasks. However, it requires non-trivial efforts to implement these methods on different models. We present LlamaFactory, a unified framework that integrates a suite of cutting-edge efficient training methods. It provides a solution for flexibly customizing the fine-tuning of 100+ LLMs without the need for coding through the built-in web UI LlamaBoard. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks. It has been released at https://github.com/hiyouga/LLaMA-Factory and received over 25,000 stars and 3,000 forks.
@@ 1 Introduction
LlamaFactory consists of three main modules: Model Loader, Data Worker and Trainer. We minimize the dependencies of these modules on specific models and datasets, allowing the framework to flexibly scale to hundreds of models and datasets. Concretely, we first establish a model registry where the Model Loader can precisely attach adapters to the pre-trained models by identifying exact layers. Then we develop a data description specification that allows the Data Worker to gather datasets by aligning corresponding columns. Furthermore, we provide plug-and-play implementations of state-of-the-art efficient fine-tuning methods that enable the Trainer to activate by replacing default ones. Our design allows these modules to be reused across different training approaches, significantly reducing the integration costs.
@@
LlamaFactory is open-sourced under the Apache-2.0 license. It has already garnered over 25,000 stars and 3,000 forks on the GitHub, and hundreds of open-source models have been built upon LlamaFactory on the Hugging Face Hub [^1] https://huggingface.co/models?other=llama-factory. For example, Truong et al. (2024) build GemSUra-7B based on LlamaFactory, revealing the cross-lingual abilities of Gemma (Mesnard et al., 2024). Furthermore, dozens of studies have utilized our framework to explore LLMs (Wang et al., 2023a; Yu et al., 2023; Bhardwaj et al., 2024).
@@ 4.1 Model Loader
This section initially presents the four components in Model Loader: model initialization, model patching, model quantization, and adapter attaching, followed by a description of our approach of adapting to a wide range of devices by handling the parameter floating-point precision during fine-tuning.
@@ 4.4 Utilities
During inference time, we reuse the chat template from the Data Worker to build the model inputs. We offer support for sampling the model outputs using Transformers (Wolf et al., 2020) and vLLM (Kwon et al., 2023), both of which support stream decoding. Additionally, we implement an OpenAI-style API that utilizes the asynchronous LLM engine and paged attention of vLLM, to provide high-throughput concurrent inference services, facilitating the deployment of fine-tuned LLMs into various applications.
@@ 4.5 LlamaBoard: A Unified Interface for LlamaFactory
LlamaBoard allows us to customize the fine-tuning arguments through interaction with the web interface. We provide default values for a majority of arguments that are recommended for most users, simplifying the configuration process. Moreover, users can preview the datasets on the web UI to validate them.
@@
LlamaBoard supports calculating the text similarity scores on the datasets to automatically evaluate models or performing human evaluation by chatting with them.